Telescope and AGw Commissioning at LFBG with IRTC - 02-03 May 2009 MST

Observers:: D.Thompson
Telescope Operator: David Gonzalez Huerta
Software Support: (phone) Telescope Support:

Summary

  • TCS problems traced to OSS, see IT#2156, 2157, 2158, 2159. Other sources (servo code)?
  • Lost 3hr to HW/SW problems
  • Collected additional pointing model data in 2-3 arcsec seeing
  • Closed at 01:25 MST to increasingly heavy clouds.

Details (Times in MST)

High cirrus at sunset. Mirror is warm so we are ventilating the dome. Afternoon tests of AGw#4 showed a problem in mcspu such that we were failing to reach "on source" (the telescope got there but the on source bits are not being correctly set. We switched back to bent Gregorian about an hour before sunset but the problem persists. Called T.Sargent who is looking into the problem. MCSPU is apparently confused as to whether it is slewing or tracking.

20:30 Tom reports it is fixed.

20:33:20.861preset to BS9143
we seem to be hung at the point where GCS is just about to ask for an acquisition image. From the GCS log:
May  2 20:36:43 lbtmu107 LBT_left_GCS: GT: setting noAzCam flag to 0
May  2 20:36:43 lbtmu107 LBT_left_GCS: GCS constructor finished.
May  2 20:36:43 lbtmu107 LBT_left_GCS: registering commands ...
May  2 20:36:49 lbtmu107 LBT_left_GCS: Registration complete
May  2 20:36:49 lbtmu107 LBT_left_GCS: starting GCS thread...
May  2 20:36:49 lbtmu107 LBT_left_GCS: GCS Subsystem Thread Started
Just so folks are not mislead, the above issue associated with GCS was actually the same problem which happened at 21:23; the OSS was hung waiting for a command to complete. Michele

20:43 It has been more than 5 minutes...shut down GCS and AGw, restart and home motors. AGW is refusing to start, giving RPC server error messages. Oacserver was not running, restarted. Restarted IIF because preset was still hanging.

20:52:12.544 preset again. We get to field and AZ, EL, and rotator appear to be tracking, but the AZtracker and ELtracker show as "slew_to_track" instead of "tracking".

21:23:09.956 preset...nope. Nothing is happening after the "wait for elevation complete" message. Michele said the OSS is hung. Restarting it immediately made the hung preset complete (with an rpc error).

22:05:00.531 preset again to BS9148 at high EL. Seems to have completed as expected (GCS took an acquisition image and identified a "star" (noise or something, see acquire000001) but the subsequent guiding failed because there was no star. See IT#2156, 2157, 2158.

Open

22:15 Opening the dome. Mirror ventilation went off at some point during this, restarting. Toutside=3.3, Tmirror=4.4 RH=65.4 wind ~10m/s with gusts to 15.

22:24:13.372 pointto ACT0224 Tel="ACQUIRE". Huge focus offset. Corrected pointing to IE=-85, CA=-60.5 (about 19 arcsec from nominal pointing model).

22:38:01.515 pointto BS9153 gs=0 Tel="ACTIVE" to collimate. WFS: 1514, 1241, 329, 356. Seeing is poor, 2-2.5 on guider.

22:43:32.570 Acquire again to update pointing. IE=-92

*Bad seeing (~2.5 arcsec) and not clear...pointing observations*
pointing log started: /home/telescope/20090503.ptlog

NN Image# Cube StarID 01 055056 0001 BS9153
00 055213 0002 BS9153sky
WT10_405 not found, getting cloudy and EL~18?
00 062454 0003 Arcturus 100x0.05 PAM-, don't use
00 062545 0004 Arcturus 100x0.0005 PAM- don't use
02 062749 0005 Arcturus 100x0.005
00 063016 0006 ACT0260 10x0.25
03 063059 0007 ACT0260 10x1.0
04 063342 0008 WT10_088 10x1.0
05 063614 0009 WT10_175 10x1.0
00 063811 0010 ACT0264 10x1.0
06 063910 0011 ACT0264 10x1.0
07 064123 0012 ACT0215
08 064318 0013 ACT0317
09 064522 0014 WT10_330
10 064650 0015 WT10_300
11 064832 0016 WT10_249
12 065118 0018 WT10_296 (0017 is 1x1.0) 13 065333 0019 ACT0192
14 065432 0020 ACT0192
15 065636 0021 WT10_077
16 065819 0022 WT10_137
00 WT10_279 hit a Y limit on M1, offset and recollimate
17 071144 0023 WT10_279
18 071407 0024 WT10_241
19 071554 0025 WT10_203
00 071732 0026 WT10_290
20 071846 0027 WT10_290
21 072125 0028 ACT0326
22 072409 0029 ACT0378
23 072629 0030 ACT0405
24 072818 0031 ACT0399
25 073013 0032 ACT0339
00 073257 0033 WT10_273 off chip
26 073500 0034 WT10_273
27 073701 0035 ACT0188
00 073911 0036 ACT0262
28 074020 0037 ACT0262
29 074349 0038 ACT0394
30 074535 0039 WT10_407
31 074708 0040 ACT0411
00 074904 0041 ACT0380 not seen in ptspiral...clouds.
32 075418 0042 WT10_409
33 075609 0043 ACT0365
34 075814 0044 WT10_372
35 080006 0045 WT10_350
36 080143 0046 ACT0352
37 080318 0047 ACT0326
00 080453 0048 WT10_311
38 080731 0049 WT10_224
39 080914 0050 WT10_152
40 081048 0051 WT10_154
41 081229 0052 ACT0209?
42 081338 0053 ACT0209
00 081606 0054 WT10_228 not visible
43 081757 0055 WT10_228
44 081927 0056 WT10_211
45 082106 0057 WT10_141
46 082247 0058 ACT0126
47 082419 0059 ACT0179 not found, very cloudy now

Close

01:25 Closing dome due to clouds.

-- JohnHill - 02 May 2009
Topic revision: r5 - 04 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback